# Federated Distributed Key Generation

### v2

@@ Abstract
Distributed Key Generation (DKG) is foundational to many threshold-based cryptographic protocols, including threshold signatures, secure multiparty computation, and i-voting systems. However, standard $(t,n)$-DKG protocols typically require a predefined set of $n$ participants and a global threshold $t$, limiting their applicability in open, decentralized settings where network membership is dynamic and node availability can be unpredictable.
We introduce Federated Distributed Key Generation (FDKG), which relaxes these constraints by allowing each participant to select its own ”guardian set” ($k$ nodes, $t$-of-$k$ local threshold) for reconstructing its partial secret key. This mechanism, inspired by Federated Byzantine Agreement, enables flexible trust delegation and enhances resilience. FDKG generalizes traditional DKG and achieves its core functionality in two communication rounds. Our simulation-based security analysis establishes correctness and privacy for generation, and liveness for reconstruction (if an adversary cannot control a participant and kt+1k-t+1 of its guardians). Simulations further confirm FDKG’s efficacy against node unreliability.
In a scenario with 100 potential parties (50 participating, $k=40$ guardians, 80% retention for tallying), FDKG distribution broadcasts total approx. $336$ kB, and reconstruction phase approx. $525$ kB. Both phases scale with $\mathcal{O}(n\cdot k)$ in terms of broadcast data. Groth16 NIZK client-side proving per participant is $\approx 5$ s for distribution, and ranges from $0.6$ s to $29.6$ s for reconstruction, depending on the number of shares processed.
Our work advances distributed cryptography by providing a robust and adaptable DKG primitive suited for dynamic networks, with potential applications ranging from ad-hoc collaborations and decentralized credential management to secure blockchain governance.
@@ I Introduction
Distributed Key Generation (DKG) protocols are essential for establishing shared cryptographic keys in decentralized systems, thereby eliminating single points of trust and failure. They form the bedrock of numerous applications, including threshold signatures, secure multi-party computation (MPC), and privacy-preserving internet voting systems where distributed trust is paramount [1]. However, classical DKG approaches often fall short in highly dynamic or permissionless environments due to their inherent assumptions about network stability and participant knowledge.
A standard $(t,n)$-DKG protocol [2] involves a fixed set of $n$ nodes collectively generating a secret key whose shares are distributed such that any $t$ or more nodes can reconstruct or use the secret, while fewer than $t$ learn nothing. This model typically requires $n$ to be known at the outset to determine polynomial degrees and sharing parameters. In open settings where participation is voluntary (e.g., public blockchains or ad-hoc networks), the actual number of participants may only become clear after an initial interaction phase. If this number deviates significantly from an anticipated $n$, a pre-selected global threshold $t$ might become unachievable [3, 4] or insecure, rendering the system ineffective. Furthermore, traditional DKG often assumes uniform trust among all $n$ participants, which may not hold in diverse decentralized ecosystems.
To overcome these limitations, we introduce Federated Distributed Key Generation (FDKG), a novel protocol that draws conceptual inspiration from Federated Byzantine Agreement (FBA) protocols [5], designed for scenarios with uncertain participation and heterogeneous trust relationships. FDKG allows any subset of potential parties to participate in the key generation process. Its core innovation lies in Guardian Sets: each participating party $P_{i}$ independently selects a trusted subset of $k$ other parties (its guardians) and defines a local threshold $t$ necessary for these guardians to reconstruct $P_{i}$’s partial secret key. This federated approach to trust delegation provides enhanced flexibility and resilience.
Formally, an FDKG protocol involves $n$ potential parties. A subset $\mathbb{D}\subseteq\mathbb{P}$ chooses to participate in generating partial secret keys. Each participant $P_{i}\in\mathbb{D}$ distributes its partial secret key $d_{i}$ among its chosen guardian set $\mathcal{G}_{i}$ of size $k$ using a local $(t,k)$-threshold SSS scheme. The global secret key d is the sum of all valid partial secrets $d_{i}$. During reconstruction, d can be recovered if, for each $P_{i}\in\mathbb{D}$, either $P_{i}$ itself provides $d_{i}$, or at least $t$ of its guardians in $\mathcal{G}_{i}$ provide their respective shares. FDKG generalizes traditional $(t,n)$-DKG, which can be viewed as a special case where all $n$ parties participate and each party’s guardian set comprises all other n1n-1 parties.
@@
  • Dynamic Participation: FDKG accommodates an arbitrary number of participants $P_{i}\in\mathbb{D}$ joining the key generation process without prior global coordination on the total number of participants.
  • Flexible Trust Model: By allowing each participant to define its own guardian set $\mathcal{G}_{i}$ and local threshold $t$, FDKG supports diverse and evolving trust relationships, making it suitable for decentralized settings where trust is not uniform. This is illustrated in Figure 1, which contrasts FDKG’s federated trust with centralized and traditional distributed trust models.
@@
In this paper, we provide a detailed specification of the FDKG protocol, a formal security analysis within the simulation-based (Real/Ideal) paradigm (Section V), and demonstrate its practical utility through an application to an internet voting scheme. Our primary contribution is the FDKG framework itself, which enables robust and verifiable key generation with adaptable trust for dynamic networks.
@@
  • We formally define the FDKG protocol, detailing its key distribution and reconstruction phases.
  • We provide a security analysis of FDKG in the simulation-based framework, proving its correctness and privacy for key generation, and analyzing the liveness conditions for key reconstruction (Section V).
@@
  • We integrate FDKG into an end-to-end internet voting scheme, showcasing its practical application using threshold ElGamal encryption and NIZK-based verifiability (Section VII). A corresponding open-source implementation is provided.
@@ II Related Work
Distributed Key Generation (DKG) is a cornerstone of threshold cryptography, enabling a set of parties to jointly compute a shared public key and individual secret key shares without any single party learning the entire secret key. Seminal DKG protocols, such as those by Gennaro et al. [2], established foundational techniques based on verifiable secret sharing (VSS) [6] over discrete logarithm-based cryptosystems, typically using Shamir’s Secret Sharing (SSS) [7] for share distribution. These protocols ensure that any authorized threshold of parties can reconstruct or use the secret, while unauthorized subsets gain no information.
Subsequent research has extensively explored DKG, focusing on various aspects such as robustness against adversarial behavior, efficiency, round complexity, and applicability in different network models (synchronous vs. asynchronous). Considerable effort has been dedicated to enhancing DKG for reliable operation in challenging environments like the internet, with a focus on asynchronous settings and high thresholds of corruptions [8, 9, 10, 11]. These works primarily address fault tolerance and adversarial resilience within a framework where the set of participants is often known or becomes known during an initial setup phase.
@@
A significant body of work focuses on Publicly Verifiable Secret Sharing (PVSS) [12, 13], where the correctness of shared information can be verified by any observer, not just the share recipients. PVSS is crucial for non-interactive DKG protocols, where parties broadcast their contributions along with proofs, minimizing interaction rounds. Groth’s work on DKG and key resharing [14] presents efficient constructions leveraging pairing-based cryptography and NIZK proofs, aiming for minimal rounds and public verifiability, which is a common goal for modern DKG deployed on transparent ledgers or bulletin boards. The security definitions and expected rigor for DKG protocols, particularly those aiming for provable security in simulation-based models, are exemplified in recent literature such as Katz’s work on round-optimal, fully secure DKG [15].
Our work on Federated Distributed Key Generation (FDKG) addresses a distinct set of challenges primarily related to dynamic participation and flexible trust delegation in settings where the total number of participants is not known beforehand, and individual participants may have heterogeneous trust in others. Traditional DKG and many NI-DKG schemes [2, 14, 15] typically assume a fixed set of $n$ participants for whom the DKG is run, or require mechanisms to establish this set. The polynomial degree in SSS, for example, depends on the intended threshold, which itself is often related to $n$. If $n$ is unknown until after the first round of messages (where parties declare participation), pre-determining a globally optimal polynomial degree becomes problematic.
Dynamic DKG schemes [16] were proposed to allow parties to join and leave an existing DKG setup. While these offer flexibility, they often involve multiple rounds of interaction for updates, which can be cumbersome for applications requiring rapid setup or involving human-in-the-loop interaction, such as ad-hoc collaborations or internet voting applications. FDKG aims to achieve its key generation and initial share distribution within a minimal number of rounds (one for distribution and one for reconstruction).
FDKG introduces participant-defined ”guardian sets,” allowing each key contributor to specify a trusted subset of parties responsible for helping reconstruct their specific partial secret key. This localized trust model, where each participant $P_{i}$ defines its own $(t,k)$-threshold for its share $d_{i}$ among $k$ guardians, contrasts with the global $(t,n)$-threshold common in standard DKG. This approach offers enhanced adaptability in unreliable or permissionless networks where global trust assumptions are hard to establish or maintain. The flexibility offered by FDKG generalizes traditional DKG protocols; a standard $(t,n)$-DKG can be seen as a special case of FDKG where every participant considers all other n1n-1 parties as their guardians.
Our design draws conceptual inspiration from Federated Byzantine Agreement (FBA) protocols, such as the Stellar Consensus Protocol [5]. FBA allows open participation by letting nodes choose their own ”quorum slices” (sets of trusted peers), enabling decentralized consensus without a globally predefined set of validators. Similarly, FDKG allows participants to define their local trust spheres (guardian sets) for key management. As a party is included in more guardian sets, its importance for liveness within the FDKG network naturally increases, reflecting a form of emergent trust or criticality.
The primary motivation for FDKG in this paper is its application to internet voting. Existing e-voting systems often rely on a pre-selected, fixed set of trusted authorities (Key Holders or Trustees) to manage decryption keys [17, 18, 19, 20, 21, 22]. While these systems use threshold cryptography to protect privacy against a subset of colluding authorities, their reliance on a static, globally trusted set makes them vulnerable if these authorities are compromised, collude beyond the threshold, or are simply unavailable during critical phases like tallying. FDKG aims to decentralize this trust by allowing individual participants (or groups acting on their behalf) to establish partial keys and manage their recovery through self-chosen guardians, enhancing system resilience and flexibility against such failures. This approach aligns with the goal of creating more robust and decentralized foundations for secure multi-party applications.

### v3

@@ Abstract
Distributed Key Generation (DKG) underpins threshold cryptography across many domains: decentralized custody and wallets, validator/committee key ceremonies, cross-chain bridges, threshold signatures, secure multiparty computation, and internet voting.
Classical $(t,n)$-DKG assumes a fixed set of $n$ parties and a global threshold $t$, and effectively requires full, timely participation; when actual participation or availability deviates, setups abort or must be rerun—a severe problem in open or ad-hoc, time-critical settings, especially when $n$ is large and node availability is unpredictable.”
We introduce Federated Distributed Key Generation (FDKG), inspired by Federated Byzantine Agreement, that makes participation optional and trust heterogeneous. Each participant selects a personal guardian set $\mathcal{G}_{i}$ of size $k$ and a local threshold $t$; its partial secret can later be reconstructed either by itself or by any $t$ of its guardians. FDKG generalizes PVSS-based DKG and completes both generation and reconstruction in a single broadcast round each, with total communication $\Theta(n\ k)$ and (in the worst case) $\mathcal{O}(n^{2})$ for reconstruction.
Our security analysis shows that: (i) Generation achieves correctness, privacy, and robustness under standard PVSS-based DKG assumptions; and (ii) Reconstruction provides liveness and privacy characterized by the guardian-set topology $\{\mathcal{G}_{i}\}$. Liveness holds provided that, for every participant $i$, the adversary does not control $i$ together with at least kt+1k-t+1 of its guardians. Conversely, reconstruction privacy holds unless the corrupted set itself is reconstruction-capable.
In a scenario with n=100n{=}100 potential parties ($|\mathbb{D}|{=}50$, k=40k{=}40, 80% retention for tallying), FDKG distribution broadcasts $\approx 336$ kB and reconstruction $\approx 525$ kB. Client-side Groth16 proving per participant is $\approx 5$ s for distribution, and 0.60.629.629.6 s for reconstruction depending on the number of revealed shares.
@@ I Introduction
Distributed Key Generation (DKG) is a multi-party protocol that lets a group of mutually distrustful parties jointly create a public/secret key pair and hold it via threshold shares: any $t$ out of $n$ parties can reconstruct or use the secret, while fewer learn nothing [1, 2]. By shifting key creation from a single entity to a committee, DKG removes single point of failure for core operations such as decryption or signing. DKGs are widely used across threshold encryption and signatures [2, 1, 3, 4, 5], consensus protocols [6], multi party computation (MPC) [7, 8, 9], distributed randomness beacons (DRB) [10, 11, 12], cryptocurrency wallets [13, 14] and internet voting protocols [15, 16, 17, 18].
However, a standard $(t,n)$-DKG protocol [1, 2] requires $n$ to be known a priori to determine polynomial degrees and sharing parameters. In settings where participation is voluntary (e.g., public blockchains or ad hoc networks), the realized participant count $n^{\prime}$ may only become clear after an initial interaction phase. If $n^{\prime}$ deviates from the anticipated $n$, the pre-selected global threshold $t$ may be unattainable ($t>n^{\prime}$) or misaligned with the actual corruption/availability assumptions; in either case the system must be reparameterized and restarted, incurring coordination cost and delay that are often unacceptable in time-critical deployments. Furthermore, traditional DKGs enforce uniform (homogeneous) trust among all $n$ parties, discarding the organic trust structure that could otherwise be leveraged to improve security and availability.
To overcome these limitations, we introduce Federated Distributed Key Generation (FDKG), a protocol that draws conceptual inspiration from Federated Byzantine Agreement (FBA) protocols [19], designed for scenarios with uncertain participation and heterogeneous trust relationships. FDKG allows any subset of potential parties to participate in the key generation process. Its core innovation lies in Guardian Sets: each participating party $P_{i}$ independently selects a trusted subset of $k$ other parties (its guardians) and defines a local threshold $t$ necessary for these guardians to reconstruct $P_{i}$’s partial secret key.
Formally, an FDKG protocol involves $n$ potential parties. A subset $\mathbb{D}\subseteq\mathbb{P}$ chooses to participate in generating partial secret keys. Each participant $P_{i}\in\mathbb{D}$ distributes its partial secret key $d_{i}$ among its chosen guardian set $\mathcal{G}_{i}$ of size $k$ using a local $(t,k)$-threshold secret sharing scheme. The global secret key d is the sum of all valid partial secrets $d_{i}$. During reconstruction, d can be recovered if, for each $P_{i}\in\mathbb{D}$, either $P_{i}$ itself provides $d_{i}$, or at least $t$ of its guardians in $\mathcal{G}_{i}$ provide their respective shares. FDKG generalizes traditional $(t,n)$-DKG, which can be viewed as a special case where all $n$ parties participate and each party’s guardian set comprises all other n1n-1 parties.
@@
  • Optional Participation: FDKG accommodates any subset of parties $P_{i}\in\mathbb{D}$ joining the key generation process without a prior commitment phase, improving UX by reducing interaction and avoiding costly reparameterizations/restarts when $n^{\prime}\neq n$.
  • Heterogeneous Trust Model: By allowing each participant to define its own guardian set $\mathcal{G}_{i}$, FDKG supports diverse and evolving trust relationships, making it suitable for decentralized settings where trust is not uniform. This is illustrated in Figure 1, which contrasts FDKG’s federated trust with centralized and traditional distributed trust models.
@@
In practice, FDKG succeeds in settings where classical $(t,n)$-DKG cannot tolerate participation uncertainty without costly restarts and timeline slips. Consider a public voting with a 24-hour enrollment window for talliers. The organizer publishes a list of $n=60$ eligible talliers and plans a classical DKG with $t=40$. After the window, only $n^{\prime}=37$ complete enrollment and verification. Since classical DKG pre-fixes $t$ against the planned $n$, the chosen $t=40$ is unattainable ($t>n^{\prime}$), so setup must abort or be restarted with new parameters. In FDKG, the same upper bound $n=60$ is published, but Round 1 accepts any subset $\mathbb{D}$ of those who post valid transcripts (here $|\mathbb{D}|=37$).
A time-critical disclosure scenario similarly benefits. A whistleblower wants to share files with up to $n=12$ independent journalists; the case is urgent, so he sets a 6-hour enrollment window. His goal is to include as many as possible by the deadline, encrypt once the window closes, and avoid being blocked by a few non-responders. With FDKG, the joint key emerges from the actual participants $\mathbb{D}$, so progress does not hinge on every individual being present. Classical DKG, by contrast, requires full participation; if fewer than the planned $n$ attend within 6 hours, the setup fails and must be rerun—missing the deadline the whistleblower cannot extend.
@@
In this paper, we provide a detailed specification of the FDKG protocol, a formal security analysis within the simulation-based paradigm (Section V), and demonstrate its practical utility through an application to an internet voting scheme. Our primary contribution is the FDKG framework itself, which enables robust and verifiable key generation with adaptable trust for dynamic networks.
@@
  • We formally define the FDKG protocol.
  • We provide a security analysis in the simulation-based framework, establishing Correctness, Privacy, and Robustness for key generation, and Liveness for reconstruction (Section V).
@@
  • We integrate FDKG into an end-to-end internet voting scheme, showcasing its practical application using threshold ElGamal encryption and NIZK-based verifiability (Section VII). A corresponding open-source implementation, together with reproducibility scripts, is provided.
@@
Table I: Comparison of synchronous, biasable DKG protocols. Core columns follow Bacho–Kavousi [20]. We split “Corrupt” into Privacy and Liveness, and add Trust and Setup. Privacy: adversary cannot learn the shared secret. Liveness: adversary cannot prevent honest parties from recovering the shared secret. Let $\mathcal{C}\subseteq\mathbb{P}$ be the statically corrupted set, and let $R$ be the family of reconstruction-capable sets (as in Eq. (1)). Communication is total broadcast cost; $BC_{n}(b)$ is the cost for one party to broadcast $b$ bits to $n$ parties; $k$ denotes the guardian set size used in the FDKG. Rounds denotes the number of broadcast rounds. Sharing indicates the secret-sharing type (VSS, PVSS, or APVSS). Trust Model: “Homo” (homogeneous) means a global, uniform trust assumption; “Hetero” (heterogeneous) allows participant-defined, non-uniform trust. Setup: “Required” means a fixed participant set must take part; “Optional” discovers participants from a larger pool.
Protocol Year Privacy Liveness Communication Rounds Sharing Trust Setup
Pedersen [1] 1991 $|\mathcal{C}|<t<n/2$ $|\mathcal{C}|<n-t$ $n\ BC_{n}(\lambda n)$ $3\cdot BC_{n}$ VSS Homo Required
Fouque–Stern [21] 2001 $|\mathcal{C}|<t<n$ $|\mathcal{C}|<n-t$ $n\ BC_{n}(\lambda n)$ $1\cdot BC_{n}$ PVSS Homo Required
Kate et al. [22] 2010 $|\mathcal{C}|<t<n/2$ $|\mathcal{C}|<n-t$ $n\ BC_{n}(\lambda+\delta\lambda n)$ $3\cdot BC_{n}$ VSS Homo Required
Gürkan et al. [23] 2021 $|\mathcal{C}|<t<\log n$ $|\mathcal{C}|<n-t$ $_{n}(\lambda)+\ell^{2}BC_{n}(\lambda n) BCn\ell\cdot BC_{n}$ APVSS Homo Required
Groth [24] 2021 $|\mathcal{C}|<t<n$ $|\mathcal{C}|<n-t$ $n\ BC_{n}(\lambda^{2}n)$ $1\cdot BC_{n}$ PVSS Homo Required
Kate et al. [25] 2023 $|\mathcal{C}|<t<n$ $|\mathcal{C}|<n-t$ $n\ BC_{n}(\lambda n)$ $1\cdot BC_{n}$ PVSS Homo Required
Cascudo–David [26] 2023 $|\mathcal{C}|<t<n$ $|\mathcal{C}|<n-t$ $n\ BC_{n}(\lambda n)$ $1\cdot BC_{n}$ PVSS Homo Required
Feng et al. [27] 2023 $|\mathcal{C}|<t<n/2$ $|\mathcal{C}|<n-t$ $\kappa\ BC_{n}(\lambda n)+n\lambda\kappa$ $2\cdot BC_{n}$ VSS Homo Required
Bacho et al. [28] 2023 $|\mathcal{C}|<t<n/2$ $|\mathcal{C}|<n-t$ $BC_{n}(\lambda n)+n^{2}\lambda\log n$ $4\cdot BC_{n}$ APVSS Homo Required
Feng et al. [29] 2024 $|\mathcal{C}|<t<n/2$ $|\mathcal{C}|<n-t$ $\sqrt{n}\ BC_{n}(\lambda n\ \kappa)$ $2\cdot BC_{n}$ PVSS Homo Required
FDKG (this work) 2024 $\boldsymbol{\mathcal{C}\notin R}$ $\boldsymbol{\exists\,S\in R:\ S\subseteq\mathbb{P}\setminus\mathcal{C}}$ $\boldsymbol{n\ BC_{n}(\lambda k)}$ $\boldsymbol{1\cdot BC_{n}}$ PVSS Hetero Optional
@@ II Related Work
Foundational DKG protocols, such as [1, 2], established the core paradigm: $n$ participants, each acting as a dealer for a partial secret, combine their contributions to form a single joint public key and distributed secret shares. These protocols are typically based on Verifiable Secret Sharing (VSS) [30], which allows participants to verify the integrity of their received shares.
@@
Subsequent research has extensively explored DKG, focusing on various aspects such as robustness against adversarial behavior, efficiency, round complexity, security, and applicability in different network models (synchronous vs. asynchronous) [20]. Considerable effort has been dedicated to enhancing DKG for reliable operation in challenging environments like the internet, with a focus on asynchronous settings and high thresholds of corruptions [31, 32, 33, 34]. A significant body of work focuses on Publicly Verifiable Secret Sharing (PVSS) [35, 36], where the correctness of shared information can be verified by any observer, not just the share recipients. PVSS is crucial for non-interactive DKG protocols, where parties broadcast their contributions along with proofs, minimizing interaction rounds. Groth’s work on DKG and key resharing [24] presents efficient constructions leveraging pairing-based cryptography and NIZK proofs, aiming for minimal rounds and public verifiability, which is a common goal for modern DKG deployed on transparent ledgers or bulletin boards.
The recent SoK by Bacho and Kavousi [20] identifies two primary axes for classifying DKG protocols: the underlying network model and the secret sharing mechanism employed. To accurately position FDKG, we adopt this framework and establish a coherent baseline for comparison.
First, DKG protocols are distinguished by their network model assumption. Synchronous protocols assume bounded message delays, allowing them to operate in well-defined rounds. In contrast, asynchronous protocols [37] are designed to tolerate unbounded message delays, a significantly stronger guarantee that requires more complex consensus mechanisms and often results in higher communication overhead. As FDKG is designed for synchronous environments, we omit asynchronous DKG protocols from our direct comparison to ensure an evaluation of protocols operating under similar network guarantees.
A second critical distinction is the protocol’s resilience to key biasing by a rushing adversary. Modern research differentiates between biasable protocols and unbiasable (or ”fully secure”) protocols [38]. Fully secure DKGs guarantee that the final public key’s distribution is uniform and cannot be influenced by a malicious minority. This robustness typically requires additional rounds of communication (e.g., commit-reveal schemes). FDKG, in line with many practical protocols, prioritizes simplicity and a minimal round count, and is therefore biasable. Consequently, to evaluate FDKG against schemes with similar design trade-offs, we focus our comparison on the class of biasable protocols.
Furthermore, it is important to distinguish DKG from Dynamic and Proactive Secret Sharing (DPSS) schemes [39]. DPSS protocols are designed for the long-term lifecycle management of a secret, enabling committees to evolve and shares to be proactively refreshed to defend against adaptive adversaries. Since FDKG focuses on the initial bootstrapping of a key, DPSS protocols address a different problem domain and are thus not directly included in our setup-focused analysis.
Our design draws conceptual inspiration from Federated Byzantine Agreement (FBA) protocols—such as the Stellar Consensus Protocol [19]—where nodes declare local “quorum slices” (trusted peers), enabling open participation without a globally pre-defined validator set. FDKG borrows this form of heterogeneous, participant-defined trust: each party $i$ specifies a guardian set $G_{i}$ with local threshold $t$, and a party included in many guardian sets becomes operationally more critical for liveness. The semantics, however, differ fundamentally. In FBA, safety and liveness are defined via quorum intersection and concern consensus over a ledger; in FDKG, properties are defined via the reconstruction family $R$ (defined in Section V) and concern a one-shot cryptographic reconstruction predicate. Thus, while both systems use local trust slices, FDKG makes no consensus claims; our guarantees (Theorems 34) address reconstruction privacy and liveness rather than agreement.
To establish a coherent baseline and facilitate a sound comparison, we therefore narrow our focus to the class of synchronous, biasable DKG protocols. Within this class, different secret sharing mechanisms are used, as detailed in the [20]. These include complaint-based VSS, non-interactive Publicly Verifiable Secret Sharing (PVSS) [40, 21] (enabling one-round designs), and Aggregatable PVSS (APVSS) [23, 41] (allowing transcript aggregation). Beyond the sharing mechanism, protocols differ in three comparison-relevant dimensions that we report in Table I: (i) the privacy security and liveness tolerance for corrupted parties; (ii) the round complexity, ranging from one broadcast round (PVSS) up to four rounds (multi-phase APVSS/complaint flows); and (iii) the communication overhead.
Our comparative analysis, presented in Table I, evaluates FDKG against this established baseline. The table reveals a consistent architectural paradigm across all surveyed protocols. To the best of our knowledge, FDKG is the first protocol in its class to break from this paradigm in two fundamental ways. First, it introduces a heterogeneous trust model, moving away from the assumption that trust is global and uniform (“Homo”). Second, it supports optional participation, whereas all prior schemes require the participation of a fixed, pre-defined set of members (“Required”). These architectural changes are what enable FDKG to provide key generation in time-critical and ad-hoc environments.